OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (aum):

First term = -10 Second term = -10 + (-2) = -12 Third term = -12 + (-2) = -14 Just keep adding -2 to get the next term.

OpenStudy (anonymous):

Thank you :)